OpenStudy (anonymous):

*help requested* *fan and medal awarded* Solve for y: 2/3y + 8 = 6 y = 3 y = 5 y = -3 y = -5 The set of ordered pairs below defines a relation. In each ordered pair (x, y), x represents the daily wages of a plumber, and y represents the daily wages of her assistant. {(144, 72), (171, 84), (198, 84), (171, 96), (144, 72)} What is the difference between the values of the least and greatest elements in the domain of this relation? 24 54 72 126

OpenStudy (anonymous):

hey are you the same girl as the pic

OpenStudy (anonymous):

because your very pretty

OpenStudy (anonymous):

yes lol but i need help

OpenStudy (anonymous):

how old are you

OpenStudy (anonymous):

im 17

OpenStudy (anonymous):

i need help. not a convo. if u want to talk msg me.

OpenStudy (matt101):

For the first question, if y is part of the denominator, I get -1/3 which isn't an option. I'm assuming it's (2/3)y + 8 = 6 meaning: (2/3)y = -2 y = -3 For your second question, the domain refers to the x coordinate in each ordered pair, so the difference between the greatest and least would be 198 - 144 = 54.

OpenStudy (anonymous):

#1C.y=-3(i know) #2C.72(i think)

OpenStudy (anonymous):

uhm well @Torrayani<3 , @matt101 did the work so ima give the medal to him. plus u got the second one wrong.